OpenStudy (anonymous):

t=-0.005x^2+0.45x+125 Will this equation ever exceed the value of 141?

jimthompson5910 (jim_thompson5910):

Plug in t = 141 and solve for x. What do you get?

OpenStudy (anonymous):

umm im not sure can u explain/

OpenStudy (anonymous):

?*

jimthompson5910 (jim_thompson5910):

t=-0.005x^2+0.45x+125 141=-0.005x^2+0.45x+125 0=-0.005x^2+0.45x+125-141 0=-0.005x^2+0.45x-16 -0.005x^2+0.45x-16 = 0 Now use the quadratic formula to solve for x Do you know the quadratic formula?

OpenStudy (anonymous):

no..... thats what im on that im struggling with...

jimthompson5910 (jim_thompson5910):

The quadratic formula is \[\Large x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] Does that look familiar?

OpenStudy (anonymous):

no.....

jimthompson5910 (jim_thompson5910):

So you haven't learned it yet?